OpenStudy (anonymous):

The Boardman family walked two over five of a mile in two over seven of an hour. What is their unit rate in miles per hour? two-sevenths over two-fifths equals one and two-fifths miles per hour two-sevenths over two-fifths equals five over seven mile per hour two-fifths over two-sevenths equals one and two-fifths miles per hour two-fifths over two-sevenths equals five over seven mile per hour

OpenStudy (anonymous):

@me11 and @angel12310

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

1. two-sevenths over two-fifths equals one and two-fifths miles per hour 2. two-sevenths over two-fifths equals five over seven mile per hour 3. two-fifths over two-sevenths equals one and two-fifths miles per hour 4. two-fifths over two-sevenths equals five over seven mile per hour

OpenStudy (anonymous):

those are the answers :D

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Vickiesweet

OpenStudy (phi):

You want miles per hour. you might have noticed that this translates to the fraction \[ \frac{miles}{hour} \] so you should (at the least) expect to divide the miles number by the hours number that rules out some of your choices

OpenStudy (anonymous):

ok

OpenStudy (phi):

which of the 4 choices shows miles divided by hours ?

OpenStudy (anonymous):

3 and 4

OpenStudy (anonymous):

right?

OpenStudy (phi):

so you have 50-50 chance if you guessed at this point. However, you can figure out the answer. remember this useful "rule" if you divide by a fraction, you can change the problem into multiply by that fraction after "flipping" it.

OpenStudy (phi):

can you "flip" 2/7 ?

OpenStudy (anonymous):

yes into 7/2

OpenStudy (phi):

now multiply 2/5 * 7/2 what do you get ?

OpenStudy (anonymous):

1.4

OpenStudy (phi):

btw, I would "cancel" the 2 because it's in the top and in the bottom

OpenStudy (phi):

I assume you used a calculator? But you can see the answer is more than 1, and only one of C and D have an answer more than 1

OpenStudy (anonymous):

no I didn't actually

OpenStudy (phi):

or change 0.4 into 4/10 = 2/5 so 1.4 is the same as 1 and 2/5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1.4 = 1 2/5

OpenStudy (phi):

yes. so now you should be able to pick out the correct choice.

OpenStudy (anonymous):

yes, its 3

OpenStudy (phi):

yes.

OpenStudy (anonymous):

thanks so much!